The width of a rectangle is 24 feet less than its length. What polynomials will represent its perimeter and area if the actual length is 44 feet?
Answer by thechamp1011(19)

TO FIND THE PERIMETER: DRAW A RECTANGLE FIRST, SINCE YOU HAVE 2 SIDES OF LENGTH AND 2 SIDES OF WIDTH.
THE PERIMETER FORMULA FOR RECTANGLE IS: P=2L+2w
SINCE WE KNOW THE THE WIDTH OF THE RECTALGLE WHICH IS 24 LESS THAN THE LENGHT, WE CAN WRITE THE WIDTH EXPRESSION AS: W=L-24
NOW SUBSTITUTE THE WIDTH EQUATION TO THE PERIMETER FORMULA TO FIND THE PERIMETER OF THE POLYNOMIALS: P=2L+2W
P=2L+(L-24)
P=3L-24
THE AREA OF THE RECTANGLE IS : A=L(W)
A=44(44-24)=44(20)
A=880FEET